how many years will it take 100g of carbon-11 to decay until there is less than 0.5g remaining if its half-life is 20 minutes?
1
Expert's answer
2020-09-23T04:36:45-0400

Solution:

The decay formula can be written as:

"N=N_0\times 2^{-\frac{t}{t_{1/2}}}"

Therefore:

"t=-log_2(\frac{N}{N_0})\times t_{1/2}=-log_2(\frac{0.5g}{100g})\times 20\,minutes\approx153\,minutes"


Answer: 153 minutes